Find without using a calculator:
(a) sin t, cos t, tan t, cot t, for t= -405 degrees

Answer by lwsshak3(11628)

Find without using a calculator:
(a) sin t, cos t, tan t, cot t, for t= -405 degrees
***
reference angle for -405˚=360-405=-45˚ (in quadrant IV)
sin t=-√2/2
cos t=√2/2
tan t=-1
cot t=-1